# Tidal perturbations of an extreme mass ratio inspiral around a Kerr black hole

### v1

@@ 1 Introduction
A central goal in black-hole perturbation theory is to obtain a tractable description of gravitational, electromagnetic, and neutrino perturbations of the Kerr spacetime. The pioneering work of Teukolsky [Teukolsky:1972my, Teukolsky:1973ha], based on the Newman–Penrose formalism [Newman:1961qr], established separable wave equations for the Weyl scalars $\psi_{0}$ and $\psi_{4}$, which encode the behavior of Kerr metric perturbations in vacuum. Together with earlier developments in Kerr geometry [Carter:1968rr, Chandrasekhar:1985kt], these results laid the mathematical foundations for the study of black-hole stability, tidal interactions, and superradiant phenomena [Press:1972zz, Teukolsky:1974yv].
@@
Working with Weyl scalars, rather than the metric perturbation itself, provides direct access to physically relevant observables such as energy and angular momentum fluxes [Chrzanowski:1975wv]. This framework includes prescriptions for quantifying the absorption of mass and angular momentum by a black hole exposed to external radiation [Poisson:2004cw]. The Teukolsky equation also enables studies of quasinormal modes and their couplings [Leaver:1985ax, Green:2022htq]. More broadly, it has been widely applied to problems ranging from the stability of the Kerr spacetime to waveform modeling, including hybrid approaches that combine black-hole perturbation theory with post-Newtonian expansions [Damour:2009wj], as well as investigations of horizon absorption and tidal torquing in Kerr spacetimes [Chatziioannou:2012gq].
@@
However, some applications require explicit knowledge of the metric perturbation rather than the Weyl scalars alone. This is the case for the computation of gravitational self-force effects, which depend on local properties of the spacetime geometry [vandeMeent:2015lxa], for second-order self-force calculations, which require the metric perturbation at first order as an input [Pound:2012nt], and more generally for extensions of black-hole perturbation theory beyond the linear regime [Loutrel:2020wbw, Ripley:2020xby].
@@
Metric reconstruction techniques based on Hertz or Debye potentials [PhysRevLett.41.203, Kegeles:1979an] allow one to recover the spacetime perturbation from the Weyl scalars. These methods are essential for the calculation of Geroch–Hansen multipole moments [Geroch:1973am, LeTiec:2020bos], the construction of initial data for binaries of rotating black holes, and the development of waveform models [Yunes:2005ve].
@@
One important application concerns the tidal deformation of black holes, which in the linear regime is described by vacuum perturbations of the Kerr metric. Previous studies address fundamental questions related to the relativistic theory of tidal Love numbers, which vanish for Schwarzschild black holes but exhibit a richer structure in the spinning case, including dissipative contributions associated with horizon absorption [Binnington:2009bb, Hartle:1973zz, Poisson:2004cw, Chatziioannou:2012gq, Chia:2020yla, LeTiec:2020bos, OSullivan:2014ywd]. The nonvanishing dissipative tidal response of Kerr black holes enters directly into the modeling of gravitational-wave signals from compact binaries: it contributes to the phase evolution and waveform amplitude, thereby affecting both parameter estimation and tests of the nature of compact objects [Tagoshi:1997jy, Alvi:2001mx, Yunes:2009ef, Yunes:2010zj, Flanagan:2007ix, Hinderer:2007mb, Baiotti:2010xh, Chatziioannou:2020pqz, Cardoso:2017cfl, Cardoso:2021qqu]. Beyond astrophysical applications, tidal perturbations also provide insight into the geometry and dynamics of black holes, for example through changes in the horizon area [Poisson:2009qj], couplings between spin and tidal fields [Poisson:2014gka], and modifications of the geodesic structure of tidally deformed spacetimes [Pani:2015hfa, Cardoso:2021qqu].
@@
We are interested in deriving the metric in the near-zone region, defined as the region at distances $r$ from the world line such that $r\ll\mathscr{R}$, where $\mathscr{R}$ denotes the tidal length scale of the external spacetime. This requirement imposes the key condition that the size of the black hole, characterized by its mass $M$, must be small compared to the tidal scale $\mathscr{R}$, i.e., $M\ll\mathscr{R}$. This hierarchy ensures that the tidal deformation can be treated perturbatively and that, on scales larger than $\mathscr{R}$, the external spacetime is not significantly affected by the presence of the black hole [Poisson:2009qj].
We apply our result for the metric perturbation of a Kerr black hole to study the secular dynamics of a test particle in the tidally deformed spacetime. This application is particularly relevant for Extreme-Mass-Ratio Inspiral (EMRI) systems, which consist of a stellar-mass compact object inspiraling into a supermassive black hole with a mass ratio of order $10^{-5}$$10^{-6}$ or smaller [Berry:2019wgg, Barack:2018yly]. EMRIs can emit gravitational-wave signals lasting up to $\mathscr{O}(10^{5})$ orbital cycles in the millihertz LISA band [Barack:2018yly], making them natural laboratories for probing secular phenomena and prime targets for space-based detectors such as LISA.
@@
Throughout this paper, we employ geometrized units with $G=c=1$ and adopt the “mostly plus” metric signature (+++)(-+++). Greek indices run over spacetime coordinates, while Latin indices run over spatial coordinates.
@@ 2 Teukolsky master equation
To compute perturbations of the Kerr metric induced by external tidal fields, we employ the Teukolsky Master Equation (TME), first introduced by Teukolsky in his pioneering works [Teukolsky:1973ha, Teukolsky:1974yv]. The TME provides a powerful framework for analyzing linear perturbations of rotating black holes in terms of gauge-invariant curvature variables.
@@
Teukolsky’s key insight was to reformulate the linearized Einstein equations in terms of curvature perturbations rather than metric perturbations. Specifically, starting from the Weyl tensor $C_{\mu\nu\rho\sigma}$ and writing the metric as in Eq. (2.1), one can extract the part of the Weyl tensor that depends linearly on the perturbation $h_{\mu\nu}$. Following Ref. [Berens:2024czo], we denote this contribution by $C_{\mu\nu\rho\sigma}^{(1)}$. Using this quantity, one can define two complex scalars, $\psi_{0}$ and $\psi_{4}$, which encode the two independent degrees of freedom of the metric perturbation $h_{\mu\nu}$. They are defined as
@@
A common choice for the Newman–Penrose tetrad is the Kinnersley tetrad $\{l^{\mu}_{\rm K},n^{\mu}_{\rm K},m^{\mu}_{\rm K},\bar{m}^{\mu}_{\rm K}\}$ [Kinnersley:1969zza], which is aligned with the principal null directions of the Kerr spacetime. In ingoing Kerr coordinates $(v,r,z,\psi)$, it takes the form [Yunes:2005ve]
@@
Even though the use of ingoing Kerr coordinates ensures the regularity of the metric at the future event horizon, this does not guarantee that the Weyl scalars $\psi_{0}$ and $\psi_{4}$ remain regular there. Regularity of the Weyl scalars can instead be achieved by adopting the Hartle–Hawking (HH) tetrad $\{l^{\mu}_{\rm HH},n^{\mu}_{\rm HH},m^{\mu}_{\rm HH},\bar{m}^{\mu}_{\rm HH}\}$ [Hawking:1972hy]. The null vectors $l^{\mu}_{\rm K}$ and $n^{\mu}_{\rm K}$ of the Kinnersley tetrad are related to those of the HH tetrad, $l^{\mu}_{\rm HH}$ and $n^{\mu}_{\rm HH}$, through the transformations
@@
while the complex null vector $m^{\mu}$ is the same in both tetrads [Poisson:2004cw, LeTiec:2020bos]. Explicitly, the HH tetrad takes the form
@@
Moreover, the Weyl scalars in the HH tetrad are related to those in the Kinnersley tetrad by [Chandrasekhar:1985kt]
@@
Here $\Psi^{(s)}$ denotes the perturbation field of spin weight $s$, and $T^{(s)}$ is the associated source term, which vanishes in vacuum. This form is obtained by projecting the Teukolsky operator given in Eq. (3.28) of Ref. [Cano:2023tmv] onto the HH tetrad. Equation (2.12) therefore applies to gravitational perturbations, corresponding to spin weights $s=\pm 2$. For $s=\pm 2$, the Weyl scalars $\psi_{0}$ and $\zeta^{4}\psi_{4}$ provide particular solutions of the TME with spin weights +2+2 and 2-2, respectively. We note that the expression of the source term $T^{(s)}$ is also tetrad dependent and therefore its explicit expression differs from that given in Ref. [Teukolsky:1972my]. However, in this work, we focus on vacuum perturbations and set the source term to zero.
@@ 2.1 Static modes
We restrict our analysis to slowly varying tidal fields, so that the TME (2.12) can be solved for static modes, i.e., with $\omega=0$. In this case, Eq. (2.14b) for the angular mode functions is satisfied by the spin-weighted spherical harmonics $_{s}Y_{\ell m}(z,\psi)$, and the separation constant reduces to [vandeMeent:2015lxa, Berens:2024czo]
@@ 3 Metric reconstruction
The reconstruction of the metric perturbation from solutions of the TME relies on the introduction of two tensorial differential operators, ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$, which, when applied to appropriate scalar potentials, yield the corresponding metric perturbation in a radiation gauge. These scalar quantities, known as Hertz potentials, satisfy an adjoint version of the TME and serve as intermediaries between the Weyl scalars $\psi_{0}$, $\psi_{4}$, and the metric perturbation $h_{\mu\nu}$ [Chrzanowski:1975wv, Wald:1973wwa, Berens:2024czo, LeTiec:2020bos].
@@
The Hertz potentials $\Psi_{\rm H}$ and $\Psi^{\prime}_{\rm H}$ are related to the Weyl scalars through fourth-order differential operators. In particular, the fields $\Psi^{(+2)}$ and $\Psi^{(-2)}$ appearing in Eqs. (3.2) and (3.1) are auxiliary Teukolsky fields of spin weight +2+2 and 2-2, respectively, and should not be confused with the Weyl scalars $\psi_{0}$ or $\zeta^{4}\psi_{4}$ [vandeMeent:2015lxa, LeTiec:2020bos, Berens:2024czo].
@@
In the Newman–Penrose formalism, the connection associated with the null tetrad is encoded in twelve complex spin coefficients $(\kappa,\rho,\sigma,\tau,\kappa^{\prime},\rho^{\prime},\sigma^{\prime},\tau^{\prime},\epsilon,\beta,\epsilon^{\prime},\beta^{\prime})$. In the GHP formalism, these coefficients enter naturally through the derivative operators, which act on quantities of GHP weight $(p,q)$ according to
@@
$\mathscr{D}=l^{\mu}\nabla_{\mu},\qquad\mathscr{D}^{\prime}=n^{\mu}\nabla_{\mu},\qquad\delta=m^{\mu}\nabla_{\mu},\qquad\delta^{\prime}=\overline{m}^{\mu}\nabla_{\mu}.$ (3.4)
In this notation, the reconstruction operators ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$ used in Eqs. (3.1) and (3.2) read [Berens:2024czo]
@@
${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\Psi$ $=-\frac{1}{2}\,n_{\mu}n_{\nu}\bigl(\eth^{\prime}-\tau^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)\Psi-\frac{1}{2}\,\overline{m}_{\mu}\overline{m}_{\nu}\bigl(\text{\thorn}^{\prime}-\rho^{\prime}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Psi$
$\quad+\frac{1}{2}\,n_{(\mu}\overline{m}_{\nu)}\Bigl[\bigl(\text{\thorn}^{\prime}-\rho^{\prime}+\overline{\rho}^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)+\bigl(\eth^{\prime}-\tau^{\prime}+\overline{\tau}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Bigr]\Psi.$ (3.5b)
@@
The remaining spin coefficients $(\kappa,\kappa^{\prime},\sigma,\sigma^{\prime})$ vanish identically in the HH tetrad. In the following, we work with perturbations in the ORG, Eq. (3.2). This choice is standard for computing physically meaningful quantities, as in this gauge the metric perturbation is transverse at future null infinity $\mathscr{I}^{+}$ and naturally isolates the outgoing gravitational radiation [Chrzanowski:1975wv, vandeMeent:2015lxa].

### v2

@@ 1 Introduction
A central goal in black-hole perturbation theory is to obtain a tractable description of gravitational, electromagnetic, and neutrino perturbations of the Kerr spacetime. The pioneering work of Teukolsky [1, 2], based on the Newman–Penrose formalism [3], established separable wave equations for the Weyl scalars $\psi_{0}$ and $\psi_{4}$, which encode the behavior of Kerr metric perturbations in vacuum. Together with earlier developments in Kerr geometry [4, 5], these results laid the mathematical foundations for the study of black-hole stability, tidal interactions, and superradiant phenomena [6, 7].
@@
Working with Weyl scalars, rather than the metric perturbation itself, provides direct access to physically relevant observables such as energy and angular momentum fluxes [8]. This framework includes prescriptions for quantifying the absorption of mass and angular momentum by a black hole exposed to external radiation [9]. The Teukolsky equation also enables studies of quasinormal modes and their couplings [10, 11]. More broadly, it has been widely applied to problems ranging from the stability of the Kerr spacetime to waveform modeling, including hybrid approaches that combine black-hole perturbation theory with post-Newtonian expansions [12], as well as investigations of horizon absorption and tidal torquing in Kerr spacetimes [13].
@@
However, some applications require explicit knowledge of the metric perturbation rather than the Weyl scalars alone. This is the case for the computation of gravitational self-force effects, which depend on local properties of the spacetime geometry [14], for second-order self-force calculations, which require the metric perturbation at first order as an input [15], and more generally for extensions of black-hole perturbation theory beyond the linear regime [16, 17].
@@
Metric reconstruction techniques based on Hertz or Debye potentials [18, 19] allow one to recover the spacetime perturbation from the Weyl scalars. These methods are essential for the calculation of Geroch–Hansen multipole moments [20, 21], the construction of initial data for binaries of rotating black holes, and the development of waveform models [22].
@@
One important application concerns the tidal deformation of black holes, which in the linear regime is described by vacuum perturbations of the Kerr metric. Previous studies address fundamental questions related to the relativistic theory of tidal Love numbers, which vanish for Schwarzschild black holes but exhibit a richer structure in the spinning case, including dissipative contributions associated with horizon absorption [23, 24, 9, 13, 25, 21, 26]. The nonvanishing dissipative tidal response of Kerr black holes enters directly into the modeling of gravitational-wave signals from compact binaries: it contributes to the phase evolution and waveform amplitude, thereby affecting both parameter estimation and tests of the nature of compact objects [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Beyond astrophysical applications, tidal perturbations also provide insight into the geometry and dynamics of black holes, for example through changes in the horizon area [37], couplings between spin and tidal fields [38], and modifications of the geodesic structure of tidally deformed spacetimes [39, 36].
@@
We are interested in deriving the metric in the near-zone region, defined as the region at distances $r$ from the world line such that $r\ll\mathscr{R}$, where $\mathscr{R}$ denotes the tidal length scale of the external spacetime. This requirement imposes the key condition that the size of the black hole, characterized by its mass $M$, must be small compared to the tidal scale $\mathscr{R}$, i.e., $M\ll\mathscr{R}$. This hierarchy ensures that the tidal deformation can be treated perturbatively and that, on scales larger than $\mathscr{R}$, the external spacetime is not significantly affected by the presence of the black hole [37].
We apply our result for the metric perturbation of a Kerr black hole to study the secular dynamics of a test particle in the tidally deformed spacetime. This application is particularly relevant for Extreme-Mass-Ratio Inspiral (EMRI) systems, which consist of a stellar-mass compact object inspiraling into a supermassive black hole with a mass ratio of order $10^{-5}$$10^{-6}$ or smaller [40, 41]. EMRIs can emit gravitational-wave signals lasting up to $\mathscr{O}(10^{5})$ orbital cycles in the millihertz LISA band [41], making them natural laboratories for probing secular phenomena and prime targets for space-based detectors such as LISA.
@@
Throughout this paper, we employ geometrized units with $G=c=1$ and adopt the “mostly plus” metric signature (,+,+,+)(-,+,+,+). Greek indices run over spacetime coordinates, while Latin indices run over spatial coordinates.
@@ 2 Teukolsky master equation
To compute perturbations of the Kerr metric induced by external tidal fields, we employ the Teukolsky Master Equation (TME), first introduced by Teukolsky in his pioneering works [2, 7]. The TME provides a powerful framework for analyzing linear perturbations of rotating black holes in terms of gauge-invariant curvature variables.
@@
Teukolsky’s key insight was to reformulate the linearized Einstein equations in terms of curvature perturbations rather than metric perturbations. Specifically, starting from the Weyl tensor $C_{\mu\nu\rho\sigma}$ and writing the metric as in Eq. (2.1), one can extract the part of the Weyl tensor that depends linearly on the perturbation $h_{\mu\nu}$. Following Ref. [42], we denote this contribution by $C_{\mu\nu\rho\sigma}^{(1)}$. Using this quantity, one can define two complex scalars, $\psi_{0}$ and $\psi_{4}$, which encode the two independent degrees of freedom of the metric perturbation $h_{\mu\nu}$. They are defined as
@@
A common choice for the Newman–Penrose tetrad is the Kinnersley tetrad $\{l^{\mu}_{\rm K},n^{\mu}_{\rm K},m^{\mu}_{\rm K},\bar{m}^{\mu}_{\rm K}\}$ [43], which is aligned with the principal null directions of the Kerr spacetime. In ingoing Kerr coordinates $(v,r,z,\psi)$, it takes the form [2]
@@
Although the Kinnersley tetrad is aligned with the principal null directions of the Kerr spacetime and has the convenient property that $l^{\mu}_{\rm K}$ is affinely parametrized, $\nabla_{l_{\rm K}}l_{\rm K}=0$, it is singular at the event horizon. This singular behavior is a property of the tetrad itself and is independent of the coordinate system used to express it. As a consequence, the associated Weyl scalars $\psi_{0}$ and $\psi_{4}$ also diverge at the horizon, despite the regularity of the Weyl tensor. To avoid this artificial singularity and ensure regularity of the curvature variables at the future event horizon, we instead adopt the Hartle–Hawking (HH) tetrad $\{l^{\mu}_{\rm HH},n^{\mu}_{\rm HH},m^{\mu}_{\rm HH},\bar{m}^{\mu}_{\rm HH}\}$ [44].
The null vectors $l^{\mu}_{\rm K}$ and $n^{\mu}_{\rm K}$ of the Kinnersley tetrad are related to those of the HH tetrad, $l^{\mu}_{\rm HH}$ and $n^{\mu}_{\rm HH}$, through the transformations
@@
while the complex null vector $m^{\mu}$ is the same in both tetrads [9, 21]. Explicitly, the HH tetrad takes the form
@@
Moreover, the Weyl scalars in the HH tetrad are related to those in the Kinnersley tetrad by [5]
@@
Here $\Psi^{(s)}$ denotes the perturbation field of spin weight $s$, and $T^{(s)}$ is the associated source term, which vanishes in vacuum. This form is obtained by projecting the Teukolsky operator given in Eq. (3.28) of Ref. [45] onto the HH tetrad. Equation (2.12) therefore applies to gravitational perturbations, corresponding to spin weights $s=\pm 2$. For $s=\pm 2$, the Weyl scalars $\psi_{0}$ and $\zeta^{4}\psi_{4}$ provide particular solutions of the TME with spin weights +2+2 and 2-2, respectively. We note that the expression of the source term $T^{(s)}$ is also tetrad dependent and therefore its explicit expression differs from that given in Ref. [1]. However, in this work, we focus on vacuum perturbations and set the source term to zero.
@@ 2.1 Static modes
We restrict our analysis to slowly varying tidal fields, so that the TME (2.12) can be solved for static modes, i.e., with $\omega=0$. In this case, Eq. (2.14b) for the angular mode functions is satisfied by the spin-weighted spherical harmonics $_{s}Y_{\ell m}(z,\psi)$, and the separation constant reduces to [14, 42]
@@ 3 Metric reconstruction
The reconstruction of the metric perturbation from solutions of the TME relies on the introduction of two tensorial differential operators, ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$, which, when applied to appropriate scalar potentials, yield the corresponding metric perturbation in a radiation gauge. These scalar quantities, known as Hertz potentials, satisfy an adjoint version of the TME and serve as intermediaries between the Weyl scalars $\psi_{0}$, $\psi_{4}$, and the metric perturbation $h_{\mu\nu}$ [8, 46, 42, 21].
@@
The Hertz potentials $\Psi_{\rm H}$ and $\Psi^{\prime}_{\rm H}$ are related to the Weyl scalars through fourth-order differential operators. In particular, the fields $\Psi^{(+2)}$ and $\Psi^{(-2)}$ appearing in Eqs. (3.2) and (3.1) are auxiliary Teukolsky fields of spin weight +2+2 and 2-2, respectively, and should not be confused with the Weyl scalars $\psi_{0}$ or $\zeta^{4}\psi_{4}$ [14, 21, 42].
@@
In the Newman–Penrose formalism, the connection associated with the null tetrad is encoded in twelve complex spin coefficients $(\kappa,\rho,\sigma,\tau,\kappa^{\prime},\rho^{\prime},\sigma^{\prime},\tau^{\prime},\epsilon,\beta,\epsilon^{\prime},\beta^{\prime})$. In the GHP formalism, these coefficients enter naturally through the derivative operators, which act on quantities of GHP weight $(p,q)$ according to [^2] We recall that a quantity $\eta$ has weights $(p,q)$ if under a frame transformation $(l,n,m)\mapsto(\lambda\bar{\lambda}l,\lambda^{-1}\bar{\lambda}^{-1}n,\lambda\bar{\lambda}^{-1}m)$ with $\lambda\in\mathbb{C}-\{0\}$ then $\eta$ transforms as $\eta\mapsto\lambda^{p}\bar{\lambda}^{q}\eta$.
@@
$\mathscr{D}=l^{\mu}\nabla_{\mu},\qquad\mathscr{D}^{\prime}=n^{\mu}\nabla_{\mu},\qquad\delta=m^{\mu}\nabla_{\mu},\qquad\delta^{\prime}=\bar{m}^{\mu}\nabla_{\mu}.$ (3.4)
In this notation, the reconstruction operators ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$ used in Eqs. (3.1) and (3.2) read [18, 19] (see also [42])
@@
${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\Psi$ $=-\frac{1}{2}\,n_{\mu}n_{\nu}\bigl(\eth^{\prime}-\tau^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)\Psi-\frac{1}{2}\,\bar{m}_{\mu}\bar{m}_{\nu}\bigl(\text{\thorn}^{\prime}-\rho^{\prime}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Psi$
$\quad+\frac{1}{2}\,n_{(\mu}\bar{m}_{\nu)}\Bigl[\bigl(\text{\thorn}^{\prime}-\rho^{\prime}+\overline{\rho}^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)+\bigl(\eth^{\prime}-\tau^{\prime}+\overline{\tau}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Bigr]\Psi.$ (3.5b)
@@
The remaining spin coefficients $(\kappa,\kappa^{\prime},\sigma,\sigma^{\prime})$ vanish identically in the HH tetrad. In the following, we work with perturbations in the ORG, Eq. (3.2). In this gauge, the metric perturbation naturally takes the form of a horizon-adapted lightcone gauge, which is particularly well-suited for describing tidal perturbations of black holes and for implementing metric reconstruction techniques based on the Hertz potential (see, e.g., Refs. [8, 9, 37, 14]).